OpenStudy (vickyisthesmartone123):

David has a hose that is 27 ft long. He wants to cut it into two pieces so that the longer piece is 3 ft more than 2 times the length of the shorter piece. How long will each piece of hose be? Shorter piece: ft Longer piece: ft

OpenStudy (vickyisthesmartone123):

@sofiaalva

OpenStudy (anonymous):

\[L=3+2S\] \[S+L=27\]

OpenStudy (anonymous):

Make to equations with 2 variables. Here L is the longer place, and S is the smaller one.

OpenStudy (vickyisthesmartone123):

Uhhhh....:\

OpenStudy (anonymous):

Do you know how to solve 2 equations with 2 variables?

OpenStudy (vickyisthesmartone123):

I don't get this that's why i'm asking you.

OpenStudy (f_jayyy):

wait hold up

OpenStudy (anonymous):

That is not right @F_Jayyy

OpenStudy (f_jayyy):

thats why i said hold up

OpenStudy (vickyisthesmartone123):

OK.....:)

OpenStudy (anonymous):

Lets solve the equations. \[L=3+2S\] \[S+L=27\]Insert eh value for L in the 2nd equation: \[S+(3+2S)=27 <=> S=8\] Insert the answer for S in one of the equations to find L. \[L=3+2*8=19\]

OpenStudy (darthvader2900):

Shorter piece: 8 Longer piece: 19

OpenStudy (darthvader2900):

Shorter piece: 8 Longer piece: 19

OpenStudy (f_jayyy):

longer piece= 19 shorter piece= 8

OpenStudy (f_jayyy):

he answered like a second before me lol

OpenStudy (vickyisthesmartone123):

lol it's ok ill medal you

OpenStudy (anonymous):

This is not a side for just getting the answers. You need to know how to get the answers / understand the work.

OpenStudy (vickyisthesmartone123):

next one!

OpenStudy (darthvader2900):

K

OpenStudy (vickyisthesmartone123):

:) I LOVE YOU

OpenStudy (vickyisthesmartone123):

JUST BECAUSE YOUR VERY NICE

OpenStudy (vickyisthesmartone123):

@F_Jayyy and

OpenStudy (darthvader2900):

Next one?